OpenStudy (anonymous):

the lengths of the sides of a triangle are y, y+ 1, and 7 centimeters. if the perimeter is 56 centimeters, what is the value of y?

OpenStudy (anonymous):

Well, the perimeter of a triangle is the sum of all three sides. We know that this will be 56 cm. We need to set up an equation to reflect the problem: Side 1 + S2 + S3 = P(erimeter) Inserting the given data: (y)+(y+1)+7=56 Add like terms: y + y = 2y 1 + 7 = 8 So, we have 2y + 8 = 56 Can you solve it from there? If not, let me know.